# P3-LLM: An Integrated NPU-PIM Accelerator for LLM Inference Using Hybrid Numerical Formats

### v2

@@ I Introduction
To satisfy the disparate demands of prefilling and decoding stages, heterogeneous LLM acceleration based on NPU and DRAM-based processing-in-memory (PIM) has been actively explored by recent literature [neupim, ianus, attacc, sk-pim, samsung-pim, pimba]. An integrated NPU-PIM accelerator leverages NPU for compute-intensive GEMM during prefilling, while PIM is suitable for memory-intensive GEMV during decoding. Despite PIM’s great promise in accelerating LLM inference, the PIM compute units (PCUs) suffer from significant area and power overhead due to the much lower transistor density in DRAM technology [upmem]. This PCU overhead severely limits the computation throughput of PIM, restricting its practical speedup mainly to single-batch inference on early-generation LLMs, whose multi-head attention and feed-forward layers exhibit an arithmetic intensity of one during decoding [llama, llama-2, gpt3]. Recently, low-batch edge LLM inference has gained much popularity [h2-llm], and state-of-the-art (SoTA) LLMs have adopted grouped-query attention (GQA) with arithmetic intensities larger than one [llama-3p1, llama-3p2, qwen-2p5, mistral]. This trend necessitates new PCU design with higher throughput, while remaining within the area constraints of DRAM technology.
One effective solution to alleviate the cost of compute units is quantization. Depending on the target quantized operands, SoTA LLM quantization algorithms can be broadly classified into three categories: (1) weight-only quantization [awq, gptq, quip-sharp, bitmod, squeeze-llm]; (2) KV-cache-only quantization [oaken, kivi, kvquant, sage-attn]; (3) weight-activation [^1] Here, we mainly refer activation quantization as quantizing both input activations and KV-cache. quantization [atom, quarot, qserve, students, mant, omniquant, ecco]. Among them, weight-only and KV-cache-only quantization are known to achieve near-lossless accuracy at 4-bit precision. However, since all other operands remain in FP16, these two methods offer limited compression and still require expensive hardware units for computation [figna, bitmod, awq]. In contrast, weight-activation quantization significantly reduces both memory footprint and computational cost, yet applying it to enable cost-effective PIM for edge LLM acceleration still poses several challenges.
@@
    Through comprehensive evaluation, we demonstrate that P3-LLM achieves higher accuracy than SoTA LLM quantization algorithms Oaken [oaken], QuaRot [quarot], and QServe [qserve], while offering an average of $4.9\times$, $2.0\times$, and $3.4\times$ speedups over the SoTA LLM accelerators HBM-PIM [samsung-pim], Ecco [ecco], and Pimba [pimba], respectively.
@@ II-B Processing In-Memory (PIM) for LLM Acceleration
DRAM-based PIM has become a promising solution to accelerate the decoding stage of LLM inference given its higher internal bandwidth tailored for memory-bound operations such as GEMV [samsung-pim, sk-pim, attacc, ianus, neupim, pimba]. Fig. 2 illustrates the architectures of two commercially available PIM devices: Samsung’s HBM-PIM [samsung-pim] and SK Hynix’s Accelerator-in-Memory (AiM) [sk-pim]. The left part of Fig. 2 shows a PIM channel consisting of 16 banks organized into 4 bank groups. One PIM compute unit (PCU) is placed near each DRAM bank to perform efficient GEMV operations by leveraging the abundant bank-level parallelism. Depending on area constraints, two banks may share the same PCU to amortize the area overhead [samsung-pim]. During LLM decoding, the DRAM bank transfers weights / KV-cache data to the PCU in 256-bit granularity (i.e., $\times16$-bit operands). Meanwhile, the input vector is sent from the host to either the PCU register file in HBM-PIM or the global buffer in AiM.
@@ II-C LLM Quantization
where $\Delta$ is the scaling factor, and $\texttt{Round}\left(\mathrm{x},\,\mathrm{Y}\right)$ is a function that rounds a value $\mathrm{x}$ to the closest value in a set $\mathrm{Y}$. This rounding process inevitably introduces error between the original and quantized operands. Numerous techniques have been proposed to reduce quantization error, such as mixed-precision quantization [gptq, awq, kivi, qserve, kvquant] and custom numerical formats [mant, bitmod, blockdialect, nvfp4, fp8, microscaling].
In mixed-precision domain, SoTA algorithms for weight-only [awq] and KV-cache-only [kivi] quantization have demonstrated near-lossless accuracy at 4-bit precision. To further alleviate the computation and memory demands of LLMs, several studies have explored weight-activation quantization [atom, qserve, quarot]. QuaRot [quarot] apply 4-bit quantization to weights, KV-cache, and activations. In contrast, QServe [qserve] adopts 4-bit quantization for weights and KV-cache, while maintaining activations at 8-bit to achieve superior accuracy.
Additionally, recent literature has explored custom numerical formats that can better adapt to the tensor distribution of LLMs [mant, bitmod, blockdialect, fp8, amove]. MANT [mant] and BlockDialect [blockdialect] develop adaptive data types that select an optimal numerical format to quantize each group of operands. However, selecting the optimal format for dynamic operands such as activations and KV-cache brings considerable runtime overhead. BitMoD [bitmod] re-purposes the redundant negative zero encoding in the 4-bit floating-point (FP4) format to minimize the weight quantization error. In addition to research efforts, custom data types have been widely adopted by industry. For instance, the flagship GPUs of NVIDIA and AMD support the 8-bit floating-point format with two variants: 4-bit exponent 3-bit mantissa (FP8-E4M3) and 5-bit exponent 2-bit mantissa (FP8-E5M2[fp8]. The latest NVIDIA Blackwell architecture introduces the NVFP4 format [nvfp4], which enhances the standard FP4 data type with a two-level scaling strategy.
@@ III Motivation
The above observations motivate our mixed-precision quantization scheme, namely W4A8KV4P8. In this scheme, weights (W) and KV-cache (KV) are quantized to 4 bits to maximize memory savings, while activations (A) and attention-scores (P) are retained at 8 bits to mitigate accuracy loss. Although attention-scores account for a small memory footprint, we argue that quantizing attention-scores further enhances hardware efficiency by allowing the $P\cdot V$ operation of the self-attention module to run on low-precision compute units. Unfortunately, the naïve W4A8KV4P8 quantization with the standard integer format results in large perplexity degradation, as highlighted in Fig. 3(b). To address this, P3-LLM proposes an operand-dependent W4A8KV4P8 quantization strategy using hybrid numerical formats (detailed in Section IV). This approach employs different numerical formats for each operand to minimize their quantization error. As a result, P3-LLM achieves better model performance than SoTA W4A8KV4 integer quantization algorithms, QuaRot [quarot] and QServe [qserve].
@@ III-B Limitations of Existing NPU-PIM Systems
Existing PIM solutions mainly focus on accelerating GEMV operations with an arithmetic intensity of one, thus facing challenges in low-batch decoding and GQA. To elucidate such limitations, we conduct roofline analysis on HBM-PIM supporting FP16 arithmetic [samsung-pim], which has $4\times$ higher bandwidth than normal HBM during PIM operations. Fig. 4 demonstrates that the performance benefits of HBM-PIM gradually disappear as the batch size (BS) and GQA group size (G) approach 4, due to limited computation throughput. On the other hand, NPU remains memory-bound even for moderate BS $\geq 16$}, making it highly desirable to increase the throughput of PIM while exploiting its abundant bandwidth. To address this, P3-LLM leverages W4A8KV4P8 quantization to enable cost-effective LLM inference on compact, low-precision hardware. Compared to HBM-PIM, P3-LLM can integrate $4\times$ PCUs under iso-compute-area constraints. Furthermore, as we will discuss in Section V-D, the low-precision PCU enables $2\times$ higher operating frequency than the FP16 PCU, which effectively doubles the peak throughput. Consequently, P3-LLM offers a superior performance roofline with $8\times$ higher throughput over HBM-PIM.
@@ III-C Limitations of Existing LLM Accelerator Co-Design
Numerous co-design solutions have been proposed to accelerate quantized LLMs, as summarized in Table I. BitMoD [bitmod] and Oaken [oaken] focus on 4-bit quantization of weights and KV-cache, respectively. But since all other operands remain in FP16, these two approaches lead to moderate memory savings and hardware efficiency. MANT [mant] and Ecco [ecco] achieve greater memory savings via W4A8KV4 quantization. Nevertheless, MANT exhibits noticeable accuracy drop under this aggressive quantization scheme since it does not address the outlier issue presented in activations and KV-cache. Whereas Ecco requires complicated online decompression of quantized data back to FP16 for computation, leading to considerable area overhead. Thus, although Ecco achieves high accuracy, its idea cannot be applied to designing low-precision PCUs for more efficient LLM decoding. On the other hand, Pimba [pimba] introduces a low-precision PIM architecture that exploits the 8-bit microscaling format [microscaling] to mitigate the area overhead of PCUs. However, it leaves the linear layer unquantized, and the FP32 accumulation pipeline in the microscaling data path limits the overall hardware efficiency gain. Unlike these proposals, P3-LLM seeks to strike a balanced trade-off among memory footprint, model accuracy, and hardware efficiency. By assigning a dedicated precision and efficient numerical format to each LLM operand, P3-LLM minimizes both quantization error and memory usage. Our careful choice of hybrid numerical formats allows to co-design low-precision PIM that significantly boosts the computation throughput under iso-compute-area constraints.
@@ IV-A KV-Cache Quantization
Profiling of KV-cache Distribution. Fig. 5 illustrates the KV-cache distribution from representative layers and heads of Llama-2-7B and Llama-3.1-8B, using Wikitext-2 dataset [wikitext] with a context length of 4K. Fig. 5(a) and (e) show that the value cache has no outlier pattern, whereas Fig. 5(b) and (f) reveal that the key cache contains distinct outlier channels, which pose significant challenges to quantization. These observations for align with previous findings that certain fixed channels in activations exhibit large outliers [smoothquant, llm-int8]. However, our profiling uncovers an additional important insight: the key cache of Llama-2-7B and Llama-3.1-8B exhibit drastically different patterns after rotary positional embedding (RoPE) is applied. In Llama-2-7B, the post-RoPE key cache has much less structured outlier patterns. In contrast, the key cache of Llama-3.1-8B shows little difference before and after RoPE. The reason is that Llama-3 supports a much longer sequence length than Llama-2 (128K vs. 4K tokens). As described in Section II-A, RoPE encodes positional information into the key cache via matrix rotation  specifically, by multiplying each key vector with a rotation matrix whose angle is proportional to the token’s position index. When the maximum supported sequence length is long (e.g., 128K in Llama-3), the rotation angles are very small for typical input context lengths (e.g., 4K) targeting edge inference scenarios [h2-llm, anda], and RoPE has minimal impact on the key-cache distribution. Given these insights, we adopt pre-RoPE and post-RoPE key-cache quantization for Llama-2 and Llama-3, respectively, to take the advantage of structured outlier channels.
@@
Quantization Format. Due to the dynamic nature of KV-cache, we employ 4-bit asymmetric integer (INT4-Asym) quantization to minimize the runtime overhead of quantization operations. Furthermore, since the value cache shows no outlier pattern and the key cache has been greatly smoothed, their uniform distributions are well-suited to the integer format.
@@ IV-B Attention-Score Quantization
In order to fully exploit the hardware efficiency of KV-cache quantization, it is essential to quantize additional operands, e.g., attention-scores, within the self-attention module. Although attention-scores have little impact on the overall memory footprint (see Section III-A), they play a critical role in the computation flow during self-attention. Consider a low-precision PCU supporting 8-bit multiplication, if the attention scores remain in FP16, the PCU cannot be fully utilized to accelerate self-attention. Instead, the quantized value cache must be transferred to NPU and rely on the FP16 compute units to perform multiplication with attention-scores, thus diminishing the advantages of high internal PIM bandwidth for LLM decoding. Unfortunately, existing KV-cache quantization approaches, such as QServe [qserve] and Oaken [oaken], typically keep attention-scores in FP16, which hinders the usage of low-precision hardware for multiplication with the value cache. A straightforward solution is to quantize attention-scores with INT8 [sage-attn], but this introduces two drawbacks. First, INT8 attention-scores cause noticeable perplexity degradation as shown in Table II. Second, as we will demonstrate in Section IV-D, INT8 can lead to significant accuracy loss for activation quantization. Since attention-scores behave the same as activations during the MAC operation on hardware, it is desirable to identify a good numerical format that balances quantization accuracy and hardware complexity.
@@ IV-C Weight Quantization
For weight quantization, we build upon an existing numerical format, BitMoD [bitmod], which adaptively remaps the redundant negative zero encoding of FP4 to some pre-defined special values. On top of the basic FP4 quantization values {± $\{\pm\,0,\,\pm\,0.5,\,\pm\,1,\,\pm\,1.5,\,\pm 2,\,\pm 3,\,\pm 4,\,\pm 6\}$, BitMoD introduces four additional special values {± $\{\pm\,5,\,\pm\,8\}$, where one of them can be selected to replace the negative zero. By searching for the optimal special value for each group of weights, it reduces the quantization error compared to asymmetric integer quantization with minimal hardware overhead. Notably, since weights and KV-cache are mapped to the same operand on the MAC hardware, a decoder is required to accommodate the data types of both weights and KV-cache (i.e., BitMoD and INT4-Asym), which we discuss in Section V-A.

### v3

@@ I Introduction
To satisfy the disparate demands of prefilling and decoding stages, heterogeneous LLM acceleration based on NPU and DRAM-based processing-in-memory (PIM) has been actively explored [neupim, ianus, attacc, samsung-pim, pimba, sk-pim, orches]. An integrated NPU-PIM accelerator leverages NPU for compute-intensive GEMM during prefilling, while PIM is suitable for memory-intensive GEMV during decoding. Despite PIM’s great promise in accelerating LLM inference, the PIM compute units (PCUs) suffer from significant area and power overhead due to the much lower transistor density in DRAM technology [upmem]. This PCU overhead severely limits the computation throughput of PIM, restricting its practical speedup mainly to single-batch inference on early-generation LLMs, whose multi-head attention and feed-forward layers exhibit an arithmetic intensity of one during decoding [llama, llama-2, gpt3]. Recently, low-batch edge LLM inference has gained much popularity [h2-llm], and state-of-the-art (SoTA) LLMs have adopted grouped-query attention (GQA) with arithmetic intensities larger than one [llama-3p1, llama-3p2, qwen-2p5, mistral]. This trend necessitates new PCU design with higher throughput, while remaining within the area constraints of DRAM technology.
One effective solution to alleviate the cost of compute units is quantization. By reducing the operand bit-width, quantization not only decreases the memory footprint, but also enables faster and more area-efficient computation on low-precision hardware. Depending on the target quantized operands, LLM quantization algorithms can be broadly classified into three categories: (1) weight-only quantization [awq, gptq, microscopiq]; (2) KV-cache-only quantization [oaken, kivi, kvquant]; (3) weight-activation [^1] Here, we mainly refer activation quantization as quantizing both input activations and KV-cache. quantization [tender, quarot, qserve, ecco, anda, mx_plus, mant]. Among them, weight-only and KV-cache-only quantization are known to achieve near-lossless accuracy at 4-bit precision. However, since all other operands remain in FP16, these two methods offer limited compression and still require expensive hardware units for computation [figna, bitmod, figlut]. In contrast, weight-activation quantization significantly reduces both memory footprint and hardware computation cost, yet applying it to enable cost-effective PIM for LLM inference still poses several challenges.
@@
    Through comprehensive evaluation, we demonstrate that P3-LLM achieves higher accuracy than SoTA LLM quantization algorithms Oaken [oaken], QuaRot [quarot], and QoQ [qserve], while offering an average of $4.9\times$, $2.0\times$, and $3.4\times$ speedups over the SoTA LLM accelerators HBM-PIM [samsung-pim], Ecco [ecco], and Pimba [pimba], respectively.
@@ II-B Processing In-Memory (PIM) for LLM Acceleration
DRAM-based PIM has become a promising solution to accelerate the decoding stage of LLM inference given its higher internal bandwidth tailored for memory-bound operations [samsung-pim, sk-pim, attacc, ianus, neupim, pimba, orches]. Fig. 2 illustrates the architectures of two commercially available PIM devices: Samsung’s HBM-PIM [samsung-pim] and SK Hynix’s Accelerator-in-Memory (AiM) [sk-pim]. The left part of Fig. 2 shows a PIM channel consisting of 16 banks organized into 4 bank groups. One PIM compute unit (PCU) is placed near each DRAM bank to perform efficient GEMV operations by leveraging the abundant bank-level parallelism. Depending on area constraints, two banks may share the same PCU to amortize the area overhead [samsung-pim]. During LLM decoding, the DRAM bank transfers weights / KV-cache data to the PCU in 256-bit granularity (i.e., $\times16$-bit operands). Meanwhile, the input vector is sent from the host to either the PCU register file in HBM-PIM or the global buffer in AiM.
@@ II-C LLM Quantization
where $\Delta$ is the scaling factor, and $\texttt{Round}\left(\mathrm{x},\,\mathrm{Y}\right)$ is a function that rounds a value $\mathrm{x}$ to the closest value in a set $\mathrm{Y}$. This rounding process inevitably introduces error between the original and quantized operands. Numerous techniques have been proposed to reduce quantization error, such as mixed-precision quantization and custom numerical formats.
In mixed-precision domain, SoTA algorithmic solutions on weight-only [awq, gptq] and KV-cache-only [kivi, kvquant] quantization have demonstrated near-lossless accuracy at 4-bit precision. To further alleviate the computation and memory demands of LLMs, several studies have explored weight-activation quantization [qserve, quarot, omniquant, students]. Meanwhile, recent literature in the architecture community has explored custom numerical formats that can better adapt to the tensor distribution of LLMs [mant, bitmod, blockdialect, fp8, microscaling, olive, mx_plus, spark, amove, microscopiq, anda]. In addition to research efforts, custom quantization formats have been widely adopted by industry. For instance, the flagship GPUs of NVIDIA and AMD support the 8-bit floating-point format with two variants: 4-bit exponent 3-bit mantissa (FP8-E4M3) and 5-bit exponent 2-bit mantissa (FP8-E5M2[fp8].
@@ III Motivation
The above observations motivate our mixed-precision quantization scheme, namely W4A8KV4P8. In this scheme, weights (W) and KV-cache (KV) are quantized to 4 bits to maximize memory savings, while activations (A) and attention-scores (P) are retained at 8 bits to mitigate accuracy loss. Although attention-scores account for a small memory footprint, we argue that quantizing attention-scores further enhances hardware efficiency by allowing the $P\cdot V$ operation of the self-attention module to run on low-precision compute units. Unfortunately, the naïve W4A8KV4P8 quantization with the standard integer format results in large perplexity degradation, as highlighted in Fig. 3(b). To address this, P3-LLM proposes an operand-dependent W4A8KV4P8 quantization strategy using hybrid numerical formats (detailed in Section IV). This approach employs dedicated numerical format for each LLM operand to minimize their quantization error. As a result, P3-LLM achieves better model performance than SoTA W4A8KV4 integer quantization algorithms, QuaRot [quarot] and QoQ [qserve].
@@ III-B Limitations of Existing NPU-PIM Systems
Existing PIM solutions mainly focus on accelerating GEMV with an arithmetic intensity of one, thus facing challenges in low-batch decoding and GQA. To elucidate such limitations, we conduct roofline analysis on HBM-PIM supporting FP16 arithmetic [samsung-pim], which has $4\times$ higher bandwidth than normal HBM during PIM operations. Fig. 4 demonstrates that the performance benefits of HBM-PIM gradually disappear as the batch size (BS) and GQA group size (G) approach 4, due to limited computation throughput. On the other hand, NPU remains memory-bound even for moderate BS $\geq 16$}, making it highly desirable to increase the compute throughput of PIM while exploiting its abundant bandwidth. To address this, P3-LLM leverages W4A8KV4P8 quantization to enable cost-effective LLM inference on compact, low-precision hardware. Compared to HBM-PIM, P3-LLM can integrate $4\times$ PCUs under iso-compute-area constraints. Furthermore, as we will discuss in Section V-D, the low-precision PCU enables $2\times$ higher operating frequency than the FP16 PCU, which effectively doubles the peak throughput. Thus, P3-LLM offers a superior roofline with $8\times$ higher throughput over HBM-PIM.
@@ III-C Limitations of Existing LLM Accelerator Co-Design
Numerous co-design solutions have been proposed by the architecture community to accelerate quantized LLMs, as summarized in Table I. BitMoD [bitmod] and Oaken [oaken] focus on 4-bit quantization of weights and KV-cache, respectively. But since all other operands remain in FP16, these two approaches lead to moderate memory savings and hardware efficiency. MANT [mant] and Ecco [ecco] achieve greater memory savings via W4A8KV4 quantization. Nevertheless, MANT exhibits noticeable accuracy drop under this aggressive quantization scheme since it does not address the outlier issue presented in activations and KV-cache. Whereas Ecco requires complicated online decompression of quantized data back to FP16 for computation, leading to considerable area overhead. Thus, although Ecco achieves high accuracy, its idea cannot be applied to designing low-precision PCUs for more efficient LLM decoding. On the other hand, Pimba [pimba] introduces a low-precision PIM architecture that exploits the 8-bit microscaling format [microscaling] to mitigate the area overhead of PCUs. However, it leaves the linear layer unquantized, and the FP32 accumulation pipeline in the microscaling data path limits the overall hardware efficiency gain. Unlike these proposals, P3-LLM seeks to strike a balanced trade-off among memory footprint, model accuracy, and hardware efficiency. By assigning a dedicated precision and efficient numerical format to each LLM operand, P3-LLM minimizes both quantization error and memory usage. Our careful choice of hybrid numerical formats allows to co-design low-precision PIM that significantly boosts the computation throughput under iso-compute-area constraints.
@@ IV-A KV-Cache Quantization
Profiling of KV-cache Distribution. Fig. 5 illustrates the KV-cache distribution from representative layers and heads of Llama-2-7B and Llama-3.1-8B, using Wikitext-2 dataset [wikitext] with a context length of 4K. Fig. 5(a, e) show that the value cache has no outlier pattern, whereas Fig. 5(b, f) reveal that the key cache contains distinct outlier channels, which poses significant challenges to quantization. These observations align with previous findings that certain fixed channels in activations exhibit large outliers [smoothquant, llm-int8]. However, our profiling uncovers an additional important insight: the key cache of Llama-2-7B and Llama-3.1-8B exhibit drastically different patterns after RoPE is applied. In Llama-2-7B, the post-RoPE key cache has much less structured outlier patterns. In contrast, the key cache of Llama-3.1-8B shows little difference before and after RoPE. The reason is that Llama-3 supports a much longer sequence length than Llama-2 (128K vs. 4K tokens). As described in Section II-A, RoPE encodes positional information into the key cache via matrix rotation  specifically, by multiplying each key vector with a rotation matrix whose angle is proportional to the token’s position index. When the maximum supported sequence length is long (e.g., 128K in Llama-3), the rotation angles are very small for typical input context lengths (e.g., 4K) targeting edge scenarios [h2-llm, anda], and RoPE has minimal impact on the key-cache distribution. Given these insights, we adopt pre-RoPE and post-RoPE key-cache quantization for Llama-2 and Llama-3, respectively, to take the advantage of structured outlier channels.
@@
Quantization Format. Due to the dynamic nature of KV-cache, we employ 4-bit asymmetric integer (INT4-Asym) quantization, which offers superior computational efficiency. Furthermore, since the value cache shows no outlier pattern and the key cache has been greatly smoothed, their uniform distributions are well-suited to the integer format.
@@ IV-B Attention-Score Quantization
In order to fully exploit the hardware efficiency of KV-cache quantization, it is essential to quantize additional operands, e.g., attention-scores, within the self-attention module. Although attention-scores have little impact on the overall memory footprint (see Section III-A), they play a critical role in the computation flow. Consider a low-precision PCU supporting 8-bit multiplication, if attention-scores remain in FP16, the PCU cannot be fully utilized to accelerate self-attention. Instead, the quantized value cache must be transferred to NPU and rely on the FP16 compute units to perform multiplication with attention-scores, thus diminishing the bandwidth advantages of PIM for LLM decoding. A straightforward solution is to quantize attention-scores with INT8 [sage-attn], but this introduces two drawbacks. First, INT8 attention-scores cause noticeable perplexity degradation as shown in Table II. Second, as we will demonstrate in Section IV-D, INT8 can lead to significant accuracy loss for activation quantization. Since attention-scores behave the same as activations during the MAC operation on hardware, it is desirable to identify a good numerical format that balances quantization accuracy and hardware complexity.
@@ IV-C Weight Quantization
For weight quantization, we build upon an existing numerical format, BitMoD [bitmod], which adaptively remaps the redundant negative zero encoding of FP4 to some pre-defined special values. On top of the basic FP4 quantization values {± $\{\pm\,0,\,\pm\,0.5,\,\pm\,1,\,\pm\,1.5,\,\pm 2,\,\pm 3,\,\pm 4,\,\pm 6\}$, BitMoD introduces four additional special values {± $\{\pm\,5,\,\pm\,8\}$, where one of them can be selected to replace the negative zero. By searching for the optimal special value for each group of weights, it reduces the quantization error compared to asymmetric integer quantization with minimal hardware overhead. Notably, since weights and KV-cache are mapped to the same operand on the MAC hardware, a decoder is required to accommodate the numerical formats of both weights and KV-cache (i.e., BitMoD and INT4-Asym), which we discuss in Section V-A.